# A Comprehensive Survey on Vector Database: Storage and Retrieval Technique, Challenge

### v2

@@ III-B2 Tree-Based Approach
where $w=p-q$ is the normal vector of the hyperplane, $x$ is any point on the hyperplane, and $b=-\frac{1}{2}(w\cdot p+w\cdot q)$ is the bias term. The formula for assigning a point x to a leaf node in a tree is:
@@
where $w_{i}$ and $b_{i}$ are the normal vector and bias term of the $i$-th split in the tree, and sign is a function that returns $1$ if the argument is positive, $-1$ if negative, and 00 if zero. The point $x$ follows the left or right branch of the tree depending on the sign of this expression, until it reaches a leaf node.
The formula for searching for the nearest neighbor of a query point $q$ in the forest is:
$\min_{x\in C(q)}d(q,x)$ (17)
@@
The advantage of Annoy is that it can uses multiple random projection trees to index the data points, which can increase the recall and robustness of the search, also reduce the memory usage and improve the speed of NNS, by creating large read-only file-based data structures that are mapped into memory so that many processes can share the same data. The performance of Annoy depends on several factors, such as the dimensionality of the data, the number of trees built, the number of nearest candidates to search, and the distance approximation method used. These factors affect the trade-off between accuracy and efficiency, as well as the complexity and scalability of the algorithm. There are also some challenges and extensions of Annoy, such as dealing with noisy and outlier data, choosing a good splitting plane for each node, or using multiple distance metrics to increase recall.
Best Bin First [80], [81]. It is a technique for finding the approximate nearest neighbors of a given vector in a large collection of vectors. It works by building a kd-tree that partitions the data points into bins, and then searching for the closest bin to the query point. The algorithm then searches within the closest bin to find the closest point to the query point. The best bin first algorithm still follows (1) (2). The advantage of best bin first is that it can reduce the search time and improve the accuracy of NNS, by focusing on the most promising bins and avoiding unnecessary comparisons with distant points. The performance of best bin first depends on several factors, such as the dimensionality of the data, the number of bins per node, the number of nearest candidates to search, and the distance approximation method used. These factors affect the trade-off between accuracy and efficiency, as well as the complexity and scalability of the algorithm. There are also some challenges and extensions of best bin first, such as dealing with noisy and outlier data, choosing a good splitting dimension and value for each node, or using multiple trees to increase recall.
K-means Tree [82]. It is a technique for clustering high-dimensional data points into a hierarchical structure, where each node represents a cluster of points. It works by applying a k-means clustering algorithm to the data points at each level of the tree, and then creating child nodes for each cluster. The process is repeated recursively until a desired depth or size of the tree is reached.
The formula for assigning a point $x$ to a cluster using the k-means algorithm is:
@@
where argmin is a function that returns the argument that minimizes the expression, and $\|\cdot\|_{2}$ denotes the Euclidean norm. The formula for assigning a point $x$ to a leaf node in a k-means tree is:
@@
where $L(x)$ is the set of leaf nodes that x belongs to, and N .center is the cluster center of node $N$. The point $x$ belongs to a leaf node if it belongs to all its ancestor nodes in the tree.
The formula for searching for the nearest neighbor of a query point $q$ in the k-means tree is:
@@
The advantage of K-means tree is that it can perform fast and accurate similarity search and retrieval of data points based on their cluster membership, by following the branches of the tree that are most likely to contain the nearest neighbors of the query point. K-means tree can also support dynamic operations, such as inserting and deleting points, by updating the tree structure accordingly. The performance of K-means tree depends on several factors, such as the dimensionality of the data, the number of clusters per node, and the distance metric used. These factors affect the trade-off between accuracy and efficiency, as well as the complexity and scalability of the algorithm. There are also some challenges and extensions of K-means tree, such as dealing with noisy and outlier points, choosing a good initialization for the k-means algorithm, or using multiple trees to increase recall.
III-B3 Graph-Based Approach
The core concept of the graph-based approach is the small-world network. A small-world network is a complex network where most nodes are not directly connected, but almost any node can be reached from another in just a few steps. This means the average path length between any two nodes is much shorter than the total number of nodes. The ”six degrees of separation” theory in sociology [83] is a concrete manifestation of a small-world network.
@@ III-B3 Graph-Based Approach
Given a query vector $q$ and the task finding the $k$ closest vectors from a set $O$, the graph-based approach uses a graph $G(V,E)$ to represent these objects, where each object $o_{i}$ corresponds to a node $v_{i}$. The small-world network graph is constructed by sequentially adding all nodes. As shown in Figure 4, when a new node is added, a list of neighboring nodes is generated using a greedy algorithm, and bidirectional connections are established between the new node and all nodes in the list. Once the graph is constructed, searching for $q$ is similar to the process of adding a new node. The search starts from a randomly selected node (with different small-world variants possibly using different selection strategies). From the list of neighbors of the current node, the node most similar to $q$ is identified. If such a node is found, it becomes the next node, and the search process is repeated. If a node has a higher similarity to $q$ than all of its neighbors, the search stops, and that node is considered the one most similar to $q$. Two types of graph-based methods are introduced: navigable small world (NSW), and hierachical navigable small world (HNSW).
Navigable Small World It is a technique that uses a graph structure to store and retrieve high-dimensional vectors based on their similarity or distance[84]. The NSW algorithm builds a graph by connecting each vector to its nearest neighbors, as well as some random long-range links that span different regions of the vector space. The idea is that these long-range links create shortcuts that allow for faster and more efficient traversal of the graph, similar to how social networks have small world properties[85].
The NSW algorithm works by using a greedy heuristic to add edges to the graph[86]. The algorithm starts with an empty graph and adds one point at a time. For each point, the algorithm finds its nearest neighbor in the graph using a random walk, and connects it with an edge. Then, the algorithm adds more edges by connecting the point to other points that are closer than its current neighbors. The algorithm repeats this process until all points are added to the graph.
The formula for finding the nearest neighbor of a point p in the graph using a random walk is:
@@
then add edge ($\text{then add edge }(p,r)$
@@
The advantage of the NSW algorithm is that it can handle arbitrary distance metrics, it can adapt to dynamic data sets, and it can achieve high accuracy and recall with low memory consumption. The NSW algorithm also uses a greedy routing strategy, which means that it always moves to the node that is closest to the query vector, until it reaches a local minimum or a predefined number of hops.
The performance of the NSW algorithm depends on several factors, such as the dimensionality of the vectors, the number of neighbors per node, the number of long-range links per node, and the number of hops per query. These factors affect the trade-off between accuracy and efficiency, as well as the complexity and scalability of the algorithm. There are also some extensions and variations of the NSW algorithm, such as hierarchical navigable small world (HNSW), which adds multiple layers of graphs, each with different scales and densities, or navigable small world with pruning (NSWP), which removes redundant links to reduce memory usage and improve search speed.
Hierachical Navigable Small World [87]. It is a state-of-the-art technique for finding the approximate nearest neighbors of a given vector in a large collection of vectors. It works by building a graph structure that connects the vectors based on their similarity or distance, and then using a greedy search strategy to traverse the graph and find the most similar vectors. The HNSW algorithm still follows (21) and (22). The HNSW algorithm also builds a hierarchical structure of the graph by assigning each point to different layers with different probabilities. The higher layers contain fewer points and edges, while the lower layers contain more points and edges. When a search query comes in, the HNSW algorithm finds the closest matching data points in the highest layer. It then proceeds layer by layer, moving downwards and finding the nearest data points in each subsequent layer based on those from the layer above. These points are considered the nearest neighbors. The algorithm continues this process in the lower layers, updating the list of nearest neighbors at each step. Once it reaches the bottom layer, the HNSW algorithm returns the data points that are closest to the search query. The algorithm uses a parameter M to control the maximum number of neighbors for each point in each layer.
The formula for assigning a point $p$ to a layer $l$ using a random probability is:
@@
where $M$ is the parameter that controls the maximum number of neighbors for each point in each layer. The algorithm assigns $p$ to layer $l$ with probability $\operatorname{Pr}[p\in l]$, and stops when it fails to assign $p$ to any higher layer.
The formula for searching for the nearest neighbor of a query point q in the hierarchical graph is:
@@
The advantage of HNSW is that it can achieve better performance than other methods of ANNS, such as tree-based or hash-based techniques. For example, it can handle arbitrary distance metrics, it can adapt to dynamic data sets, and it can achieve high accuracy and recall with low memory consumption. HNSW also uses a hierarchical structure that allows for fast and accurate search, by starting from the highest layer and moving down to the lowest layer, using the closest node as the next hop at each layer. So, HNSW can effectively “jump over” large portions of data that don’t require searching. Suppose a data store has only a single layer. In that case, the search algorithm is unable to bypass unrelated objects, meaning it has to examine many more data points, even if they are unlikely to be relevant matches.
The performance of HNSW depends on several factors, such as the dimensionality of the vectors, the number of layers, the number of neighbors per node, and the number of hops per query. These factors affect the trade-off between accuracy and efficiency, as well as the complexity and scalability of the algorithm. There are also some challenges and extensions of HNSW, such as finding the optimal parameters for the graph construction, dealing with noise and outliers, and scaling to very large data sets. Some of the extensions include optimized product quantization (OPQ), which combines HNSW with product quantization to reduce quantization distortions, or product quantization network (PQN), which uses a neural network to learn an end-to-end product quantizer from data.
III-B4 Quantization-Based Approach
The core idea of quantization is to map points in a high-dimensional space to a low-precision representation in a finite set. This reduces the number of bits required for storage, lowering storage demands. By using these quantized representations during queries, the distance between the original points can be quickly estimated. Specifically, a vector quantizer maps k-dimensional vectors from the vector space $R^{k}$ to a finite set of vectors $S=\{s_{i}:i=1,\dots,n\}$. Each vector $s_{i}$ is called a code vector or codeword or centroids. The collection of all codewords is referred to as a codebook. Associated with each codeword, $s_{i}$, is a nearest neighbor region called Voronoi region, and it is defined by: $V_{i}=\left\{x\in R^{k}:\left\|x-y_{i}\right\|\leq\left\|x-y_{j}\right\|,\text{ forall }j\neq i\right\}$. The set of Voronoi region partitions the entire space $R^{k}$ such that:
$\bigcup_{i=1}^{N}V_{i}=R^{k}$
$\bigcap_{i=1}^{N}V_{i}=\phi\text{ for all }i\neq j$
Figure 5 shows the codewords in a two-dimensional space, where input vectors are marked with blue stars, codewords are marked with orange circles, and the Voronoi regions are separated with boundary lines.
@@ III-B4 Quantization-Based Approach
A vector quantizer consists of two primary components: an encoder and a decoder, as shown in figure6. The encoder takes an input vector and outputs the index of the codeword that minimizes the distortion. This minimal distortion is found by calculating the distance between the input vector and each codeword in the codebook, typically using metrics like Euclidean or Hamming distance. Once the codeword with the smallest distance is identified, its index is transmitted to the decoder. At the receiver, the decoder then maps this index back to the corresponding codeword. Generating an effective codebook involves selecting codewords that best represent a given set of input vectors, along with determining the appropriate number of codewords. Designing an optimal codebook is an NP-hard problem, implying that finding the absolute best set of codewords through exhaustive search becomes impractically complex as the number of codewords increases. As a result, heuristic methods, such as the Linde-Buzo-Gray (LBG) algorithm, which is conceptually similar to the K-means clustering algorithm, are commonly employed. To create a codebook using the LBG algorithm, one first specifies the number of codewords, $N$, which defines the size of the codebook. Initially, $N$ codewords are selected at random, often from the set of input vectors themselves. Each input vector is then associated with the nearest codeword based on the Euclidean distance. After all vectors have been assigned to their respective clusters, a new set of codewords is generated by computing the average of the vectors within each cluster. This process involves summing the components of the vectors in each cluster and dividing by the total number of vectors in that cluster. This iterative procedure refines the codebook until a satisfactory level of representation is achieved. The formula for calculating the average of the components within each cluster is:
$y_{i}=\frac{1}{m}\sum_{j=1}^{m}x_{ij}$
where $i$ is the component of each vector , $m$ is the number of vectors in the cluster. There are also many other methods to designing the codebook, methods such as Generative Pre-trained Transformer Vector quantization (GPTVQ)[88], Vector PostTraining Quantization (VPTQ)[89], deep network architecture for vector quantization (DeepVQ)[90], etc.
@@
Based on the core idea described above, six quantization-based methods will be presented here, namely inverted file index (IVF), product quantization (PQ)[91, 92], optimized product quantization (OPQ)[93, 94], online product quantization[95], scalable nearest neighbor (ScaNN)[96], and inverted file product quantization (IVF_PQ)[40, 97]. Product quantization can reduce the memory footprint and search time of ANN search, by comparing codes instead of the original vectors  [98].
Inverted File Index. Inverted File Index is a technique designed to enhance search efficiency by narrowing the search area through the use of neighbor partitions or clusters[99]. It uses clustering (e.g., K-means) to partition high-dimensional vectors into multiple regions (Voronoi Cells) and records the vectors within each region through an inverted index. During a query, the search is restricted to a few regions closest to the query vector, significantly reducing the search space and improving retrieval efficiency. IVF is often combined with other techniques, such as Product Quantization (PQ), to further optimize storage and computation, making it widely used in image retrieval, recommendation systems, and VDBs. Its main advantages are fast search speed and high efficiency, though its performance in high-dimensional spaces may be limited by clustering quality and the complexity of dynamic updates.
Product Quantization . It is a technique for compressing high-dimensional vectors into smaller and more efficient representations[91, 92]. It works by dividing a vector into several sub-vectors, and then applying a clustering algorithm (such as k-means) to each sub-vector to assign it to one of a finite number of possible values (called centroids). The result is a compact code that consists of the indices of the centroids for each sub-vector.
The PQ algorithm works by using a vector quantization technique to map each subvector to its nearest centroid in a predefined codebook. The algorithm first splits each vector into $m$ equal-sized subvectors, where m is a parameter that controls the length of the code. Then, for each subvector, the algorithm learns $k$ centroids using the k-means algorithm, where $k$ is a parameter that controls the size of the codebook. Finally, the algorithm assigns each subvector to its nearest centroid and concatenates the centroid indices to form the code.
The formula for splitting a vector $x$ into $m$ subvectors is:
@@
where $x_{i}$ is the $i$-th subvector of $x$, and has dimension $d/m$, where $d$ is the dimension of $x$.
The formula for finding the centroids of a set of subvectors $P$ using the k-means algorithm is:
@@
where $c_{i}$ is the $i$-th centroid, $S_{i}$ is the set of subvectors assigned to the i -th cluster, and ||$\cdot$|| ⋅ | denotes the cardinality of a set.
The formula for assigning a subvector $x$ to a centroid using the k-means algorithm is:
@@
where argmin is a function that returns the argument that minimizes the expression, and $\|\cdot\|_{2}$ denotes the Euclidean norm.
The formula for encoding a vector $x$ using PQ is:
@@
The advantage of product quantization is that it is simple and easy to implement, as it only requires a standard clustering algorithm and a simple distance approximation method.
The performance of product quantization depends on several factors, such as the dimensionality of the data, the number of sub-vectors, the number of centroids per sub-vector, and the distance approximation method used. These factors affect the trade-off between accuracy and efficiency, as well as the complexity and scalability of the algorithm. There are also some challenges and extensions of product quantization, such as dealing with noisy and outlier data, optimizing the space decomposition and the codebooks, or adapting to dynamic data sets.
Optimized Product Quantization [93]. It is a variation of product quantization (PQ), which is a technique for compressing high-dimensional vectors into smaller and more efficient representations. OPQ works by optimizing the space decomposition and the codebooks to minimize quantization distortions. OPQ can improve the performance of PQ by reducing the loss of information and increasing the discriminability of the codes [94].
The advantage of OPQ is that it can achieve higher accuracy and recall than PQ, as it can better preserve the similarity or distance between the original vectors.
The formula for applying a random rotation to the data is:
@@
where $P_{i}$ is the set of subvectors assigned to the $i$-th cluster, $R_{i}$ is the rotation matrix for the $i$-th cluster, and $c_{i}$ is the quantization function for the $i$-th cluster, which returns the nearest centroid in the codebook.
The formula for encoding a vector $x$ using OPQ is:
@@
where $x_{i}$ is the $i$-th subvector of $x$, $R_{i}$ is the rotation matrix for the $i$th subvector, and $q_{i}$ is the quantization function for the $i$-th subvector, which returns the index of the nearest centroid in the codebook.
The performance of OPQ depends on several factors, such as the dimensionality of the data, the number of sub-vectors, the number of centroids per sub-vector, and the distance approximation method used. These factors affect the trade-off between accuracy and efficiency, as well as the complexity and scalability of the algorithm. There are also some challenges and extensions of OPQ, such as dealing with noisy and outlier data, choosing a good optimization algorithm, or combining OPQ with other techniques such as hierarchical navigable small world (HNSW) or product quantization network (PQN).
Online Product Quantization [100]. It is a variation of product quantization (PQ), which is a technique for compressing high-dimensional vectors into smaller and more efficient representations. Online product quantization (O-PQ) works by adapting to dynamic data sets, by updating the quantization codebook and the codes online. O-PQ can handle data streams and incremental data sets, without requiring offline retraining or reindexing.
The formula for splitting a vector $x$ into $m$ subvectors is:
@@
where $x_{i}$ is the $i$-th subvector of $x$, and has dimension $d/m$, where $d$ is the dimension of $x$.
The formula for initializing the centroids of a set of subvectors $P$ using the k-means++ algorithm is:
@@
where $c_{i}$ is the $i$-th centroid, with probability proportional to $D(x)^{2}$, $D(x)$ is the distance between point $x$ and its closest centroid among $\left\{c_{1},\ldots,c_{i-1}\right\}$.
The formula for assigning a subvector $x$ to a centroid using $\mathrm{PQ}$ is:
@@
where $\arg\min$ is a function that returns the argument that minimizes the expression, and $\|\cdot\|_{2}$ denotes the Euclidean norm.
The formula for encoding a vector $x$ using $\mathrm{PQ}$ is:

### v3

@@ 3.2.2 Tree-Based Approach
where w=pqw=p-q is the normal vector of the hyperplane, $x$ is any point on the hyperplane, and $b=-\frac{1}{2}(w\cdot p+w\cdot q)$ is the bias term. The formula for assigning a point x to a leaf node in a tree is:
@@
where $w_{i}$ and $b_{i}$ are the normal vector and bias term of the $i$-th split in the tree, and sign is a function that returns $1$ if the argument is positive, 1-1 if negative, and 0 if zero. The point $x$ follows the left or right branch of the tree depending on the sign of this expression, until it reaches a leaf node. The formula for searching for the nearest neighbor of a query point $q$ in the forest is:
$\vskip-5.69054pt\min_{x\in C(q)}d(q,x)$ (17)
@@
The advantage of Annoy is that it can use multiple random projection trees to index the data points, which can increase the recall and robustness of the search, also reduce the memory usage, and improve the speed of NNS by creating large read-only file-based data structures that are mapped into memory so that many processes can share the same data. The performance of Annoy depends on several factors, such as the dimensionality of the data, the number of trees built, the number of nearest candidates to search, and the distance approximation method used.
Best Bin First [beis1997shape, liu2011improved]. It is a technique for finding the approximate nearest neighbors of a given vector in a large collection of vectors. It works by building a kd-tree that partitions the data points into bins and then searching for the closest bin to the query point. The algorithm then searches within the closest bin to find the closest point to the query point. The best bin first algorithm still follows (1)  (2). The advantage of best bin first is that it can reduce the search time and improve the accuracy of NNS, by focusing on the most promising bins and avoiding unnecessary comparisons with distant points. The performance of the best bin first depends on several factors, such as the dimensionality of the data, the number of bins per node, the number of nearest candidates to search, and the distance approximation method used.
K-means Tree [tavallali2021k]. It is a technique for clustering high-dimensional data points into a hierarchical structure, where each node represents a cluster of points. It works by applying a k-means clustering algorithm to the data points at each level of the tree and then creating child nodes for each cluster. The process is repeated recursively until a desired depth or size of the tree is reached. The formula for assigning a point $x$ to a cluster using the k-means algorithm is:
@@
where argmin is a function that returns the argument that minimizes the expression and $\|\cdot\|_{2}$ denotes the Euclidean norm. The formula for assigning a point $x$ to a leaf node in a k-means tree is:
@@
where $L(x)$ is the set of leaf nodes that $x$ belongs to, and $N.center$ is the cluster center of node $N$. The point $x$ belongs to a leaf node if it belongs to all its ancestor nodes in the tree. The formula for searching for the nearest neighbor of a query point $q$ in the k-means tree is:
@@
The advantage of K-means tree is that it can perform fast and accurate similarity searches and retrieval of data points based on their cluster membership by following the branches of the tree that are most likely to contain the nearest neighbors of the query point. K-means tree can also support dynamic operations, such as inserting and deleting points, by updating the tree structure accordingly. The performance of K-means tree depends on several factors, such as the dimensionality of the data, the number of clusters per node, and the distance metric used.
3.2.3 Graph-Based Approach
The core concept of the graph-based approach is the small-world network. A small-world network is a complex network where most nodes are not directly connected, but almost any node can be reached from another in just a few steps. This means the average path length between any two nodes is much shorter than the total number of nodes. The "six degrees of separation" theory in sociology [guare2016six] is a concrete manifestation of a small-world network.
@@ 3.2.3 Graph-Based Approach
Given a query vector $q$ and the task of finding the $k$ closest vectors from a set $O$, the graph-based approach uses a graph $G(V,E)$ to represent these objects, where each object $o_{i}$ corresponds to a node $v_{i}$. The small-world network graph is constructed by sequentially adding all nodes. As shown in Figure 4, when a new node is added, a list of neighboring nodes is generated using a greedy algorithm, and bidirectional connections are established between the new node and all nodes in the list. Once the graph is constructed, searching for $q$ is similar to the process of adding a new node. The search starts from a randomly selected node (with different small-world variants possibly using different selection strategies). From the list of neighbors of the current node, the node most similar to $q$ is identified. If such a node is found, it becomes the next node, and the search process is repeated. If a node has a higher similarity to $q$ than all of its neighbors, the search stops, and that node is considered the one most similar to $q$. Two types of graph-based methods are introduced: navigable small world (NSW) and hierarchical navigable small world (HNSW).
Navigable Small World It is a technique that uses a graph structure to store and retrieve high-dimensional vectors based on their similarity or distance [ponomarenko2011approximate]. The NSW algorithm builds a graph by connecting each vector to its nearest neighbors, as well as some random long-range links that span different regions of the vector space. The idea is that these long-range links create shortcuts that allow for faster and more efficient traversal of the graph, similar to how social networks have small world properties [malkov2014approximate].
The NSW algorithm works by using a greedy heuristic to add edges to the graph[malkov2012scalable]. The algorithm starts with an empty graph and adds one point at a time. For each point, the algorithm finds its nearest neighbor in the graph using a random walk and connects it with an edge. Then, the algorithm adds more edges by connecting the point to other points that are closer than its current neighbors. The algorithm repeats this process until all points are added to the graph. The formula for finding the nearest neighbor of a point p in the graph using a random walk is:
@@
$\text{then add edge }(p,r)$
@@
The advantage of the NSW algorithm is that it can handle arbitrary distance metrics, it can adapt to dynamic data sets, and it can achieve high accuracy and recall with low memory consumption. The NSW algorithm also uses a greedy routing strategy, which means that it always moves to the node that is closest to the query vector until it reaches a local minimum or a predefined number of hops. The performance of the NSW algorithm depends on several factors, such as the dimensionality of the vectors, the number of neighbors per node, the number of long-range links per node, and the number of hops per query.
Hierachical Navigable Small World [malkov2018efficient]. It is a state-of-the-art technique for finding the approximate nearest neighbors of a given vector in a large collection of vectors. It works by building a graph structure that connects the vectors based on their similarity or distance and then using a greedy search strategy to traverse the graph and find the most similar vectors. The HNSW algorithm still follows (21) and (22). The HNSW algorithm also builds a hierarchical structure of the graph by assigning each point to different layers with different probabilities. The higher layers contain fewer points and edges, while the lower layers contain more points and edges. When a search query comes in, the HNSW algorithm finds the closest matching data points in the highest layer. It then proceeds layer by layer, moving downwards and finding the nearest data points in each subsequent layer based on those from the layer above. These points are considered the nearest neighbors. The algorithm continues this process in the lower layers, updating the list of nearest neighbors at each step. Once it reaches the bottom layer, the HNSW algorithm returns the data points that are closest to the search query. The algorithm uses a parameter M to control the maximum number of neighbors for each point in each layer. The formula for assigning a point $p$ to a layer $l$ using a random probability is:
@@
where $M$ is the parameter that controls the maximum number of neighbors for each point in each layer. The algorithm assigns $p$ to layer $l$ with probability $\operatorname{Pr}[p\in l]$, and stops when it fails to assign $p$ to any higher layer. The formula for searching for the nearest neighbor of a query point q in the hierarchical graph is:
@@
The advantage of HNSW is that it can achieve better performance than other methods of ANNS, such as tree-based or hash-based techniques. For example, it can handle arbitrary distance metrics, it can adapt to dynamic data sets, and it can achieve high accuracy and recall with low memory consumption. The performance of HNSW depends on several factors, such as the dimensionality of the vectors, the number of layers, the number of neighbors per node, and the number of hops per query.
Beyond the greedy traversal inherent to graph-based methods, further acceleration can be achieved through geometric pruning techniques that leverage metric space properties. Triangle inequalities, for example, can safely eliminate nodes unlikely to be closer than the current best candidate during graph search, reducing distance computations. This principle extends beyond graphs to other index structures. Recent work demonstrates this by applying fine-grained triangle inequality pruning within cluster-based indexes, achieving over 99.4% pruning ratio[xu2025tribase]. Such geometric pruning strategies are orthogonal to the underlying index structure—whether graph, tree, or cluster-based—and can be integrated for compounded performance gains.
3.2.4 Quantization-Based Approach
The core idea of quantization is to map points in a high-dimensional space to a low-precision representation in a finite set. This reduces the number of bits required for storage, lowering storage demands. By using these quantized representations during queries, the distance between the original points can be quickly estimated. Specifically, a vector quantizer maps k-dimensional vectors from the vector space $R^{k}$ to a finite set of vectors $S=\{s_{i}:i=1,\dots,n\}$. Each vector $s_{i}$ is called a code vector or codeword, or centroids. The collection of all codewords is referred to as a codebook. Associated with each codeword, $s_{i}$, is a nearest neighbor region called Voronoi region, and it is defined by: $V_{i}=\left\{x\in R^{k}:\left\|x-y_{i}\right\|\leq\left\|x-y_{j}\right\|,\text{ forall }j\neq i\right\}$. The set of Voronoi region partitions the entire space $R^{k}$ such that: $\bigcup_{i=1}^{N}V_{i}=R^{k}$, $\bigcap_{i=1}^{N}V_{i}=\phi\text{ for all }i\neq j$. Figure 5 shows the codewords in a two-dimensional space, where input vectors are marked with blue stars, codewords are marked with orange circles, and the Voronoi regions are separated with boundary lines.
@@ 3.2.4 Quantization-Based Approach
A vector quantizer consists of two primary components: an encoder and a decoder, as shown in Figure 6. The encoder takes an input vector and outputs the index of the codeword that minimizes the distortion. This minimal distortion is found by calculating the distance between the input vector and each codeword in the codebook, typically using metrics like Euclidean or Hamming distance. Once the codeword with the smallest distance is identified, its index is transmitted to the decoder. At the receiver, the decoder then maps this index back to the corresponding codeword. Generating an effective codebook involves selecting codewords that best represent a given set of input vectors, along with determining the appropriate number of codewords. Designing an optimal codebook is an NP-hard problem, implying that finding the absolute best set of codewords through exhaustive search becomes impractically complex as the number of codewords increases. As a result, heuristic methods, such as the Linde-Buzo-Gray (LBG) algorithm, which is conceptually similar to the K-means clustering algorithm, are commonly employed. To create a codebook using the LBG algorithm, one first specifies the number of codewords, $N$, which defines the size of the codebook. Initially, $N$ codewords are selected at random, often from the set of input vectors themselves. Each input vector is then associated with the nearest codeword based on the Euclidean distance. After all vectors have been assigned to their respective clusters, a new set of codewords is generated by computing the average of the vectors within each cluster. This process involves summing the components of the vectors in each cluster and dividing by the total number of vectors in that cluster. This iterative procedure refines the codebook until a satisfactory level of representation is achieved. The formula for calculating the average of the components within each cluster is: $y_{i}=\frac{1}{m}\sum_{j=1}^{m}x_{ij}$, where $i$ is the component of each vector and $m$ is the number of vectors in the cluster. There are also many other methods for designing the codebook, methods such as Generative Pre-trained Transformer Vector quantization (GPTVQ) [van2024gptvq], Vector PostTraining Quantization (VPTQ) [liu2024vptq], deep network architecture for vector quantization (DeepVQ) [le2018deepvq], etc.
@@
Based on the core idea described above, six quantization-based methods will be presented here, namely inverted file index (IVF), product quantization (PQ) [jegou2010product, matsui2018survey], optimized product quantization (OPQ) [ge2013optimized, li2020optimized], online product quantization[noauthor_online_nodate], scalable nearest neighbor (ScaNN) [guo2020accelerating], and inverted file product quantization (IVF_PQ) [guo2022manu, wang2021milvus]. Product quantization can reduce the memory footprint and search time of ANN search by comparing codes instead of the original vectors  [8653810]. More recently, RaBitQ [gao2024rabitq], a randomized quantization method that provides strong theoretical guarantees on distance estimation error while maintaining practical efficiency. These approaches are fundamental to modern high-dimensional similarity search.
Inverted File Index. Inverted File Index is a technique designed to enhance search efficiency by narrowing the search area through the use of neighbor partitions or clusters [liu2022new]. It uses clustering (e.g., K-means) to partition high-dimensional vectors into multiple regions (Voronoi Cells) and records the vectors within each region through an inverted index. During a query, the search is restricted to a few regions closest to the query vector, significantly reducing the search space and improving retrieval efficiency. IVF is often combined with other techniques, such as Product Quantization (PQ), to further optimize storage and computation, making it widely used in image retrieval, recommendation systems, and VDBs. Its main advantages are fast search speed and high efficiency, though its performance in high-dimensional spaces may be limited by clustering quality and the complexity of dynamic updates.
Product Quantization. It is a technique for compressing high-dimensional vectors into smaller and more efficient representations [jegou2010product, matsui2018survey]. It works by dividing a vector into several sub-vectors, and then applying a clustering algorithm (such as k-means) to each sub-vector to assign it to one of a finite number of possible values (called centroids). The result is a compact code that consists of the indices of the centroids for each sub-vector.
The PQ algorithm works by using a vector quantization technique to map each subvector to its nearest centroid in a predefined codebook. The algorithm first splits each vector into $m$ equal-sized subvectors, where $m$ is a parameter that controls the length of the code. Then, for each subvector, the algorithm learns $k$ centroids using the k-means algorithm, where $k$ is a parameter that controls the size of the codebook. Finally, the algorithm assigns each subvector to its nearest centroid and concatenates the centroid indices to form the code. The formula for splitting a vector $x$ into $m$ subvectors is:
@@
where $x_{i}$ is the $i$-th subvector of $x$, and has dimension $d/m$, where $d$ is the dimension of $x$. The formula for finding the centroids of a set of subvectors $P$ using the k-means algorithm is:
@@
where $c_{i}$ is the $i$-th centroid, $S_{i}$ is the set of subvectors assigned to the i-th cluster, and $|\cdot|$ denotes the cardinality of a set. The formula for assigning a subvector $x$ to a centroid using the k-means algorithm is:
@@
where argmin is a function that returns the argument that minimizes the expression and $\|\cdot\|_{2}$ denotes the Euclidean norm. The formula for encoding a vector $x$ using PQ is:
@@
The advantage of product quantization is that it is simple and easy to implement, as it only requires a standard clustering algorithm and a simple distance approximation method. The performance of product quantization depends on several factors, such as the dimensionality of the data, the number of sub-vectors, the number of centroids per sub-vector, and the distance approximation method used.
Optimized Product Quantization  [ge2013optimized]. It is a variation of product quantization (PQ), which is a technique for compressing high-dimensional vectors into smaller and more efficient representations. OPQ works by optimizing the space decomposition and the codebooks to minimize quantization distortions. OPQ can improve the performance of PQ by reducing the loss of information and increasing the discriminability of the codes  [li2020optimized]. The advantage of OPQ is that it can achieve higher accuracy and recall than PQ, as it can better preserve the similarity or distance between the original vectors. The formula for applying a random rotation to the data is:
@@
where $P_{i}$ is the set of subvectors assigned to the $i$-th cluster, $R_{i}$ is the rotation matrix for the $i$-th cluster, and $c_{i}$ is the quantization function for the $i$-th cluster, which returns the nearest centroid in the codebook. The formula for encoding a vector $x$ using OPQ is:
@@
where $x_{i}$ is the $i$-th subvector of $x$, $R_{i}$ is the rotation matrix for the $i$ -th subvector, and $q_{i}$ is the quantization function for the $i$-th subvector, which returns the index of the nearest centroid in the codebook.
The performance of OPQ depends on several factors, such as the dimensionality of the data, the number of sub-vectors, the number of centroids per sub-vector, and the distance approximation method used.
Online Product Quantization  [xu2018online]. It is a variation of product quantization (PQ), which is a technique for compressing high-dimensional vectors into smaller and more efficient representations. Online product quantization (O-PQ) works by adapting to dynamic data sets by updating the quantization codebook and the codes online. O-PQ can handle data streams and incremental data sets without requiring offline retraining or reindexing. The formula for splitting a vector $x$ into $m$ subvectors is:
@@
where $x_{i}$ is the $i$-th subvector of $x$, and has dimension $d/m$, where $d$ is the dimension of $x$. The formula for initializing the centroids of a set of subvectors $P$ using the k-means++ algorithm is:
@@
where $c_{i}$ is the $i$-th centroid, with probability proportional to $^{2}, D(x)D(x)$ is the distance between point $x$ and its closest centroid among $\left\{c_{1},\ldots,c_{i-1}\right\}$. The formula for assigning a subvector $x$ to a centroid using $\mathrm{PQ}$ is:
@@
where $\arg\min$ is a function that returns the argument that minimizes the expression and $\|\cdot\|_{2}$ denotes the Euclidean norm. The formula for encoding a vector $x$ using $\mathrm{PQ}$ is: